# LaST0: Latent Spatio-Temporal Chain-of-Thought for Robotic Vision–Language–Action Model

### v1

@@ I Introduction
Recently, robot imitation learning has emerged as a dominant paradigm for acquiring manipulation policies from expert demonstrations [25, 16, 72]. By inheriting the semantic understanding and common-sense reasoning capabilities of Vision–Language Models (VLMs) [1, 40, 75, 21], Vision–Language–Action (VLA) models integrate rich pretrained knowledge with the low-level control capabilities of robotic policies [6, 5, 50, 35]. This integration endows robotic agents with a unified framework for interpreting human instructions and executing corresponding manipulation primitives in dynamic environments.
@@
Rather than simply mapping observations to actions, recent advances in VLA models have been inspired by the Chain-of-Thought (CoT) reasoning paradigm in general VLMs [30, 89]. In this line of work, some approaches enhance manipulation stability and interpretability by explicitly generating linguistic reasoning traces or affordance representations [85, 46]. In parallel, other studies seek to capture environmental dynamics by predicting future states, such as subgoal images or camera-view depth maps [71, 90, 77, 92, 52]. Despite their demonstrated benefits, explicit CoT VLA methods remain constrained by two fundamental challenges in robotic manipulation. On the one hand, explicit reasoning typically incurs non-negligible inference latency. The autoregressive generation paradigm introduces inevitable computational overhead [69, 76], limiting the VLA model’s ability to achieve real-time responsiveness. This latency further restricts VLA models from reasoning effectively along the temporal dimension, thereby undermining the temporal consistency required for closed-loop manipulation. On the other hand, explicit reasoning is often confined to the linguistic space, imposing a representational bottleneck that struggles to faithfully capture ineffable physical attributes. In contrast, robotic agents must reason about and interact with the physical world, which is essential for robust manipulation in a dynamic environment.
@@
In this paper, we propose LaST0, a dual-system VLA model that enables efficient reason-before-act behavior through a Latent Spatio-Temporal Chain-of-Thought (CoT). As shown in Fig. LABEL:fig:intro, unlike prior explicit CoT-based VLA methods, LaST0 performs reasoning in a compact latent space, enabling the capture of fine-grained physical and robotic dynamics that are difficult to verbalize, while supporting temporally coherent modeling for closed-loop manipulation. Specifically, we introduce a token-efficient latent CoT space that autoregressively predicts future latent tokens of 2D images, 3D point clouds, and proprioceptive states. As a result, the VLA model can implicitly model the semantic and geometric structure of physical dynamics, while forming an internal representation of the robot state, thereby capturing the relationship between the robot and its interactive environment. Meanwhile, the latent CoT space is extended across future keyframes, enabling temporally consistent causal reasoning, which improves action coherence in closed-loop robotic manipulation. Although the proposed latent CoT is compact and encodes richer physical information, incorporating it into action generation still introduces additional inference overhead. Therefore, leveraging temporally extended latent conditions, we further propose a dual-system architecture implemented via a Mixture-of-Transformers (MoT) design, as shown in Fig. LABEL:fig:intro. Specifically, two experts are integrated within a single VLA model: a slow reasoning expert, which performs low-frequency latent inference to capture spatio-temporal dependencies, and a fast acting expert, which generates actions conditioned on high-frequency observations and periodically updated latent representations. Through shared self-attention mechanisms, LaST0 enables long-context interaction between the latent CoT space and the action space, thereby effectively coordinating deliberative reasoning with responsive control.
@@
For the training procedure, both the latent reasoning expert and the action expert are initialized from the same pretrained multimodal foundation model (i.e., Janus-Pro [12]). We then perform large-scale pretraining of LaST0 on diverse robotic manipulation datasets [62, 41, 82], ensuring a shared representation space and seamless interaction between the two experts within a unified VLA model. During downstream training, we first optimize the slow reasoning expert via supervised fine-tuning using ground-truth latent representations extracted from auxiliary encoders. This latent-space teacher forcing establishes a reliable latent scaffold for subsequent action learning. Afterward, the reasoning expert is frozen, and the action expert is trained under heterogeneous fast–slow operating ratios, enabling the VLA model to adaptively select different frequencies during deployment. For evaluation, we systematically assess LaST0 on 10 simulated tasks and six complex real-world tasks, including single-arm and dual-arm settings. Across these benchmarks, LaST0 outperforms prior state-of-the-art (SOTA) VLA methods by 8% and 13%, respectively, while achieving a 14× speedup over previous explicit CoT VLA approaches. In addition, we validate the closed-loop manipulation capability of LaST0 on long-horizon real-world tasks under dynamic conditions. Notably, the robot is able to repeatedly and continuously scoop eggs from a pan while adapting to dynamic environmental changes, demonstrating robust spatio-temporal understanding. Our contributions are summarized as follows:
@@
  • We design a spatio-temporal latent CoT space for manipulation, which autoregressively models future semantic, geometric, and proprioceptive information, allowing LaST0 to reason about physical dynamics in a temporally coherent manner.
@@
II Related Work
II-A Vision-Language-Action Model
@@
Vision-Language-Action (VLA) models have rapidly advanced as an effective paradigm for instruction-conditioned robotic manipulation, primarily driven by scaling robot demonstration data and adapting pretrained VLMs for robotic control [7, 6, 3, 48, 42]. To improve expressivity for continuous actions, recent VLA research has increasingly employed continuous generative policy heads. Diffusion-based VLA [51, 79, 81, 50] models complex action distributions through iterative denoising, while flow-matching formulations [5, 35, 22, 4, 80] offer an alternative that can improve sampling efficiency and stability. These directions aim to leverage the generalization of VLMs to specifically enhance the controllability, precision, and fidelity of continuous action generation. Beyond imitation learning, more work explores post-training VLA models via Reinforcement Learning (RL) to enhance robustness. Offline methods leverage preference-based alignment [91] to steer policies using static data without requiring simulation. Online RL methods [70, 15, 14, 44, 54] refine policies through active environment interaction. However, scaling RL-based strategies for open-world manipulation remains constrained by challenges in sample efficiency and training stability. Meanwhile, recent research equips VLA models with “reason-before-act” components to retain accurate representations of the dynamic environment and improve action quality.  [35, 49, 86] adopts textual chain-of-thought (CoT) generation for future task planning and action generation. Subsequent work [78, 92, 10, 26, 8, 90] extends generative text planning to future image prediction. Furthermore, in order to model fine-grained future states, [52, 9, 28] introduces predictions of future multi-modal information. However, explicit reasoning typically incurs non-negligible inference latency, thereby limiting the VLA model’s ability to achieve real-time responsiveness. Unlike explicit CoT-based VLA methods, LaST0 employs a latent spatio-temporal CoT strategy to efficiently capture fine-grained physical and robotic dynamics within a compact latent space.
II-B Latent Chain-of-Thought
@@
Recent work of VLM in the general domain [13, 23, 31, 84, 76] has explored latent CoT reasoning to address the limitations of explicit CoT on ineffable visual-spatial matching and high-cost text or image generation. These methods perform multi-step inference directly in continuous latent spaces, allowing intermediate reasoning to be compact, implicit, and tightly integrated with downstream prediction. Beyond general-purpose VLMs, similar approaches have been adopted in the embodied intelligence domain, where textual CoT is particularly ill-suited for representing low-level signal generation. LCDrive [69] replaces language-based explanations with action-aligned latent rollouts, allowing the model to simulate future outcomes and evaluate alternatives directly in the action space. Meanwhile, Thinkact [34] compresses intermediate motion plans into compact representations to guide action generation. Prior latent CoT frameworks mainly operate on textual reasoning or abstract 2D representations for general tasks. In contrast, LaST0 is tailored for robotic manipulation, where the VLA model reasons in a physically grounded latent space that jointly encodes semantic intent, geometric structure, and robot state, thereby capturing the embodied interaction between the robot and its environment.
III Method
@@
Figure 2: Overall framework of LaST0. a) We propose LaST0, a unified VLA model with a dual-system architecture. The model is implemented via a MoT scheme with two experts interacting through shared self-attention. The slow reasoning expert operates at a low frequency, taking visual observations and text as input to construct the LaST CoT by autoregressively predicting spatio-temporal latent tokens, which are stored in the KV cache. The fast acting expert operates at a higher frequency and generates actions via flow matching, conditioned on high-frequency observations and periodically updated latent representations. b) We design a spatio-temporal latent space tailored for the reasoning expert, where pretrained modality-specific encoders extract features from future RGB images, point clouds, and robot states at keyframes. These features serve as ground-truth latent CoT targets for supervising the reasoning expert. c) The training procedure consists of three stages with different parameter update strategies, ensuring a reliable latent and robust action generation.
In Section III-A, we describe the preliminaries of VLA models and contrast explicit and latent Chain-of-Thought (CoT) paradigms in the robotic domain. The proposed LaST0 with a Mixture-of-Transformers (MoT) architecture, featuring dual experts coupled via a shared attention mechanism, is detailed in Section III-B. Section III-C elucidates the construction of our spatio-temporal latent space, explaining the encoding and compression of future physical knowledge. We then describe the fast–slow system design and its asynchronous inference frequencies in Section III-D. Finally, Section III-E outlines the complete training paradigm, from large-scale pretraining to supervised fine-tuning.
III-A Preliminaries
@@
VLA Problem Formulation. We formulate the robot manipulation task as a probabilistic sequence decision-making problem [42]. At each timestep $t$, the policy receives a natural language instruction $l_{t}$ and visual observations $I_{t}\in\mathbb{R}^{H\times W\times 3}$ that capture the current environment. The objective of the VLA model $\pi_{\theta}$ is to generate an optimal action sequence $\mathbf{a}_{t:t+H}$ conditioned on the instruction $l_{t}$. Consistent with prior works[50, 5], we define the action space within the Special Euclidean group $SE(3)$. For single-arm configurations (e.g., Franka research 3), we employ a 7-DoF end-effector pose control mechanism, formulated as $\mathbf{a}_{t}\in\mathbb{R}^{7}$. Specifically, this control vector consists of 3-DoF for relative positional offsets ($\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DoF for rotation (represented as Euler angles [roll, pitch, yaw] $\in\mathbb{R}^{3}$), and 1-DoF for the gripper state (open/closed, $g\in\mathbb{R}^{1}$). For dual-arm configurations, to validate our model’s scalability, we extend this to a 14-DoF representation by concatenating the control signals.
Existing methods typically adopt an explicit CoT paradigm, where $\mathcal{Z}$ consists of discrete natural language tokens [45] or image tokens [92, 88]. While interpretable, these approaches confine reasoning to the linguistic space of LLMs, which struggles to represent ineffable physical attributes and incurs inference latency due to explicit decoding. To overcome these limitations, we propose a new latent CoT paradigm for the robotic domain. Unlike explicit CoT, we define $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k}\}$ as a sequence of continuous embeddings in a high-dimensional latent space. In our framework, the latent variable $\mathcal{Z}$ is trained to autoregressively predict future dynamics, including latent representations of 2D images, 3D point clouds, and robot proprioceptive states, thereby modeling the physical world in a compact space.
III-B LaST0 Architecture
@@
Overview. As illustrated in Fig. 2 a), we elaborate on the architectural of LaST0. Our model is initialized on the Janus-Pro[12], utilizing DeepSeek-LLM 1B as the backbone. To bridge the gap between reasoning and control, we transform this standard decoder-only transformer into a unified MoT dual-system architecture, constructing the ”reason-before-act” paradigm. This design enables the model to effectively decouple the generation of slow, high-level latent reasoning from fast, low-level action execution, while maintaining seamless information flow through a shared attention mechanism.
@@
Vision Encoder. For each input RGB observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$ ($H=W=224$), we employ SigLIP-Large[87] to extract semantic features. Following standard practices, the image is divided into non-overlapping patches of size $14\times 14$, which are then flattened and linearly projected. This process yields a compact feature sequence $f_{\text{img}}\in\mathbb{R}^{B\times N_{\text{img}}\times d_{v}}$, where $B$ denotes the batch size, $N_{\text{img}}$ represents the sequence length, and $d_{v}$ is the embedding dimension. In our framework, these encoded features $f_{\text{img}}$ serve a dual purpose: the current frame acts as real-time contextual input to the MoT experts, while future frames provide ground-truth target embeddings for the visual component of the latent CoT.
@@
Point Cloud Encoder (training only). To equip the model with fine-grained geometric perception and robust 3D spatial reasoning, we integrate a preyrained point cloud encoder that explicitly captures object geometry and spatial knowlegde. We adopt Uni3D [94], a scalable 3D foundation model, which has been pre-trained on massive 3D datasets to align point cloud features with image and text embeddings. Note that, unlike the vision encoder, the point cloud encoder is not used during inference. Instead, Uni3D is solely employed to encode ground-truth point clouds into compact 3D feature representations within the latent CoT space.
@@
Mixture-of-Transformers LLM Backbone. We adopt DeepSeek-LLM 1B as our foundation, repurposing its 24-layer decoder-only transformer architecture into a unified dual-system policy. Inspired by [20], to efficiently support both high-level reasoning and low-level execution, we transform the standard backbone into a MoT architecture. Unlike conventional transformers that apply a homogeneous set of weights to all tokens, our MoT design introduces task-specific parameter sets for all non-embedding components, including Feed-Forward Networks (FFN), attention projections ($W_{Q},W_{K},W_{V},W_{O}$), and Layer Normalizations, while maintaining a shared global self-attention context. This modification essentially yields two specialized experts residing within the same $d=2048$ dimensional latent space, while a slow reasoning expert responsible for autoregressively synthesizing the Latent CoT embeddings $\mathcal{Z}$ from language and slow-stream visual inputs, and a fast acting expert dedicated to generating high-frequency actions $\mathbf{a}_{t}$. The design of different operating frequencies for the two experts is described in Section III-D.
@@
MLP Components. To further clarify the LaST0 architecture, we describe the remaining auxiliary components, all of which are implemented as Multi-Layer Perceptrons (MLPs). First, a 3D Projector aligns point cloud features to the same dimension, serving as supervise targets for the Latent CoT. Regarding the action generation mechanism, given that our fast acting expert adopts a Flow Matching policy [5], specific modules are incorporated to handle the continuous generative dynamics. For the action expert input, we use a timestep MLP to encode the continuous time coordinate $t\in[0,1]$, initialized with sinusoidal embeddings, and a noised-action MLP to project the perturbed action state. For the action expert output, a projector MLP is used to transform the predicted flow velocity field, enabling the model to generate precise actions.
III-C Latent Spatio-Temporal Chain-of-Thought

### v3

# LaST0: Latent Spatio-Temporal Chain-of-Thought for Robotic Vision-Language-Action Model

@@ Abstract
Vision-Language-Action (VLA) models have recently shown strong generalization, with some approaches seeking to explicitly generate linguistic reasoning traces or predict future observations prior to execution. However, explicit reasoning typically incurs non-negligible inference latency, which constrains the temporal resolution required for robotic manipulation. Moreover, such reasoning is confined to the linguistic space, imposing a representational bottleneck that struggles to faithfully capture ineffable physical attributes. To mitigate these limitations, we propose LaST0, a framework that enables efficient reasoning before acting through a Latent Spatio-Temporal Chain-of-Thought (CoT), capturing fine-grained physical and robotic dynamics that are often difficult to verbalize. Specifically, we introduce a token-efficient latent CoT space that models future visual dynamics, 3D structural information, and robot proprioceptive states, and further extends these representations across time to enable temporally consistent implicit reasoning trajectories. Furthermore, LaST0 adopts a dual-system architecture implemented via a Mixture-of-Transformers design, where a reasoning expert conducts low-frequency latent inference and an acting expert generates high-frequency actions conditioned on robotics-oriented latent representations. To facilitate coordination, LaST0 is trained with heterogeneous operation frequencies, enabling adaptive switching during deployment. Across 10 real-world tasks spanning tabletop, mobile, and dexterous hand manipulation, LaST0 improves mean success rates by 13%, 14% and 14% over prior SOTA VLA methods, respectively.
Zhuoyang Liu $^{*\text{ }1,2,4}$, Jiaming Liu  $^{*\text{ }\dagger\text{ }1,4}$, Hao Chen $^{*\text{ }3,4}$, Jiale Yu1, Ziyu Guo3, Chengkai Hou1,2, Chenyang Gu1,
Xiangju Mi1, Renrui Zhang3, Kun Wu2, Zhengping Che $^{\dagger\text{ }2}$, Jian Tang2, Pheng-Ann Heng3, Shanghang Zhang🖂 $^{\text{\Letter}\text{ }1}$
1State Key Laboratory of Multimedia Information Processing, School of Computer Science, Peking University
2Beijing Innovation Center of Humanoid Robotics    3CUHK    4Simplexity Robotics
@@ 1 Introduction
By inheriting the semantic understanding and common-sense reasoning capabilities of Vision-Language Models (VLMs) (Alayrac et al., 2022; Karamcheti et al., 2024; Deng et al., 2025b), Vision-Language-Action (VLA) models integrate rich pretrained knowledge with the low-level control capabilities of robotic policies (Brohan et al., 2023; Black et al., 2024; Intelligence et al., 2025). This integration endows robotic agents with a unified framework for interpreting human instructions and executing corresponding manipulation primitives in dynamic environments.
@@
Rather than simply mapping observations to actions, recent advances in VLA models have been inspired by the Chain-of-Thought (CoT) reasoning paradigm in general VLMs (Guo et al., 2025). In this line of work, some approaches enhance manipulation stability and interpretability by explicitly generating linguistic reasoning traces or affordance representations (Ye et al., 2025; Li et al., 2025a). In parallel, other studies seek to capture environmental dynamics by predicting future states (Tian et al., 2024; Zhang et al., 2025b; Wang et al., 2025c; Zhao et al., 2025). Despite their demonstrated benefits, explicit CoT VLA methods remain constrained by two fundamental challenges in robotic manipulation. On the one hand, explicit reasoning typically incurs non-negligible inference latency. The autoregressive generation paradigm introduces inevitable computational overhead (Tan et al., 2025; Wang et al., 2025b), limiting the VLA model’s ability to achieve real-time responsiveness. This latency further restricts VLA models from reasoning effectively along the temporal dimension, thereby undermining the temporal consistency required for closed-loop manipulation. On the other hand, explicit reasoning is often confined to the linguistic space, imposing a representational bottleneck that struggles to faithfully capture ineffable physical attributes. In contrast, robotic agents must reason about and interact with the physical world, which is essential for robust manipulation in a dynamic environment.
@@
In this paper, we propose LaST0, a dual-system VLA model that enables efficient reason-before-act behavior through a Latent Spatio-Temporal Chain-of-Thought (CoT). As shown in Fig. LABEL:fig:intro, unlike prior explicit CoT-based VLA methods, LaST0 performs reasoning in a compact latent space, enabling the capture of fine-grained physical and robotic dynamics that are difficult to verbalize, while supporting temporally coherent modeling. Specifically, we introduce a token-efficient latent CoT space that autoregressively predicts future latent tokens of 2D images, 3D point clouds, and proprioceptive states. As a result, the VLA model can implicitly model the semantic and geometric structure of physical dynamics, while forming an internal representation of the robot state, thereby capturing the relationship between the robot and its interactive environment. Meanwhile, the latent CoT space is extended across future keyframes, enabling temporally consistent causal reasoning, which improves action coherence in closed-loop robotic manipulation. Although the proposed latent CoT is compact and encodes richer physical information, incorporating it into action generation still introduces additional inference overhead. Therefore, leveraging temporally extended latent conditions, we further propose a dual-system architecture implemented via a Mixture-of-Transformers (MoT) design. Specifically, two experts are integrated within a single VLA model: a slow reasoning expert, which performs low-frequency latent inference to capture spatio-temporal dependencies, and a fast acting expert, which generates actions conditioned on high-frequency observations and periodically updated latent representations. Through shared self-attention mechanisms, LaST0 enables long-context interaction between the latent CoT space and the action space, thereby effectively coordinating deliberative reasoning with responsive control.
@@
For the training procedure, both the latent reasoning expert and the action expert are initialized from the same pretrained VLM (i.e., Janus-Pro (Chen et al., 2025b)). We then perform large-scale pretraining of LaST0 on diverse robotic manipulation datasets (Open X-Embodiment Collaboration et al., 2023; Khazatsky et al., 2024; Wu et al., 2025; Hou et al., 2025), ensuring seamless interaction between the two experts within a unified VLA model. During downstream training, we jointly optimize the two experts, with the action expert trained under heterogeneous fast-slow operating ratios, enabling the VLA model to adaptively select appropriate execution frequencies during deployment. For evaluation, we systematically assess LaST0 on 10 simulated tasks and 10 complex real-world tasks spanning tabletop single- and dual-arm, mobile, and dexterous hand manipulation, where it outperforms prior SOTA VLA methods by 8% in simulation and by 13%, 14%, and 14% across three real-world scenarios, respectively, while achieving a 14× speedup over previous explicit CoT VLA approaches. In addition, we validate the manipulation capability of LaST0 on long-horizon real-world tasks, such as repeatedly scooping eggs from a pan while adapting to dynamic environmental changes. Our contributions are summarized as follows:
@@
  • We design a spatio-temporal latent CoT space, which autoregressively models future semantic, geometric, and proprioceptive information, allowing LaST0 to reason about physical dynamics in a temporally coherent manner.
@@
2 Related Work
@@
Vision-Language-Action (VLA) Model. VLA models are primarily driven by scaling robot demonstration data and adapting pretrained VLMs for robotic control (Belkhale et al., 2024; Li et al., 2023; Kim et al., 2024). To improve expressivity for continuous actions, recent VLA research has increasingly employed continuous generative policy heads. Diffusion-based VLA (Wen et al., 2024, 2025b; Liu et al., 2025a) models complex action distributions through iterative denoising, while flow-matching formulations (Intelligence et al., 2025; Deng et al., 2025c; Bjorck et al., 2025; Wen et al., 2025a) offer an alternative that can improve sampling efficiency and stability. Meanwhile, recent research equips VLA models with “reason-before-act” components to improve physical world reasoning.  (Intelligence et al., 2025; Lin et al., 2025; Zawalski et al., 2025) adopts textual chain-of-thought (CoT) generation for future task planning and action generation. Subsequent work (Wang et al., 2025d; Zhao et al., 2025; Chen et al., 2025a; Gao et al., 2025; Bu et al., 2025; Zhang et al., 2025b) extends generative text planning to future image prediction. Furthermore,  (Liu et al., 2025b; Cen et al., 2025; Gu et al., 2025) introduces predictions of future multi-modal information. However, explicit reasoning typically incurs non-negligible inference latency. LaST0 proposes a latent CoT strategy to efficiently capture fine-grained physical and robotic dynamics within a compact latent space.
@@
Latent CoT. Recent work of VLM in the general domain (Chen et al., 2025c; Deng et al., 2024; Hao et al., 2024; Yang et al., 2025; Wang et al., 2025b; Li et al., 2025b) has explored latent CoT reasoning to address the limitations of explicit CoT on ineffable visual-spatial matching and high-cost generation. These methods perform multi-step inference directly in continuous latent spaces, allowing intermediate reasoning to be compact, implicit, and tightly integrated with downstream prediction. Beyond general-purpose VLMs, similar approaches have been adopted in the embodied intelligence domain, where textual CoT is particularly ill-suited for representing low-level signal generation. LCDrive (Tan et al., 2025) replaces language-based explanations with action-aligned latent rollouts for autonomous driving. Thinkact (Huang et al., 2025) compresses intermediate motion plans into compact representations. Our proposed LaST0 is tailored for robotic manipulation, reasoning in a physically grounded latent space that jointly encodes semantic intent, geometric structure, and robot state.
3 Method
@@
Figure 2: Framework. a) We propose LaST0, a unified VLA model with a dual-system architecture. The model is implemented via a MoT scheme with two experts interacting through shared self-attention. The slow reasoning expert operates at a low frequency, taking images and text as input to construct the LaST CoT through autoregression. The fast acting expert operates at a higher frequency and generates actions via flow matching, conditioned on high-frequency observations and periodically updated latent representations. b) We design a spatio-temporal latent space, where pretrained modality-specific encoders extract features from future RGB images, point clouds, and robot states. These features serve as ground-truth latent CoT targets for supervising the reasoning expert.
3.1 Preliminaries
@@
VLA Problem Formulation. We formulate the robot manipulation task as a probabilistic sequence decision-making problem (Kim et al., 2024). At each timestep $t$, the policy receives a natural language instruction $l_{t}$ and visual observations $I_{t}\in\mathbb{R}^{H\times W\times 3}$ that capture the current environment. The objective of the VLA model $\pi_{\theta}$ is to generate an optimal action sequence $\mathbf{a}_{t:t+H}$ conditioned on the instruction $l_{t}$. We define the action space within the Special Euclidean group $SE(3)$. For single-arm configurations (e.g., Franka research 3), we employ a 7-DoF end-effector pose control mechanism, formulated as $\mathbf{a}_{t}\in\mathbb{R}^{7}$. Specifically, this control vector consists of 3-DoF for relative positional offsets ($\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DoF for rotation (represented as Euler angles [roll, pitch, yaw] $\in\mathbb{R}^{3}$), and 1-DoF for the gripper state (open/closed, $g\in\mathbb{R}^{1}$). For dual-arm configurations, we extend the action representation to 14 DoF by concatenating control signals to validate scalability; for mobile manipulation, we additionally estimate the base’s linear and angular velocities. Additional robotic CoT reasoning preliminaries are provided in Appendix F.1.
3.2 LaST0 Architecture
@@
Overview. As illustrated in Fig. 2 a), we elaborate on the architectural of LaST0. Our model is initialized on the Janus-Pro(Chen et al., 2025b), utilizing DeepSeek-LLM 1.5B as the backbone. To bridge the gap between reasoning and control, we transform this standard decoder-only transformer into a unified MoT dual-system architecture, constructing the ”reason-before-act” paradigm. This design enables the model to effectively decouple the generation of slow, high-level latent reasoning from fast, low-level action execution, while maintaining seamless information flow through a shared attention mechanism.
@@
Vision Encoder. For each input RGB observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$ ($H=W=384$), we employ SigLIP-Large(Zhai et al., 2023) to extract semantic features. The encoder yields a compact feature sequence $f_{\text{img}}\in\mathbb{R}^{B\times N_{\text{img}}\times d_{v}}$, where $B$ denotes the batch size, $N_{\text{img}}$ represents the sequence length, and $d_{v}$ is the embedding dimension. In our framework, these encoded features $f_{\text{img}}$ serve a dual purpose: the current frame acts as real-time contextual input to the MoT experts, while future frames provide ground-truth target embeddings for the visual component of the latent CoT.
@@
Point Cloud Encoder (training only). To equip the model with fine-grained geometric perception and robust 3D spatial reasoning, we integrate a large-scale, pretrained point cloud encoder, Uni3D (Zhou et al., 2023b), which explicitly captures object geometry and spatial knowledge. Note that, unlike the vision encoder, the point cloud encoder is not used during inference. Instead, Uni3D is solely employed to encode ground-truth point clouds into compact 3D feature representations within the latent CoT space.
@@
Mixture-of-Transformers LLM Backbone. We adopt DeepSeek-LLM 1.5B as our foundation, repurposing its 24-layer decoder-only transformer architecture into a unified dual-system policy. To efficiently support both high-level reasoning and low-level execution, we transform the standard backbone into a MoT architecture (Deng et al., 2025a). Unlike conventional transformers that apply a homogeneous set of weights to all tokens, our MoT design introduces task-specific parameter sets for all non-embedding components, including Feed-Forward Networks (FFN), attention projections ($W_{Q},W_{K},W_{V},W_{O}$), and Layer Normalizations, while maintaining a shared global self-attention context. This modification essentially yields two specialized experts residing within the same $d=2048$ dimensional latent space, while a slow reasoning expert responsible for autoregressively synthesizing the Latent CoT embeddings $\mathcal{Z}$ from language and slow-stream visual inputs, and a fast acting expert dedicated to generating high-frequency actions $\mathbf{a}_{t}$. The design of different operating frequencies for the two experts is described in Section 3.4.
@@
MLP Components. To further clarify the LaST0 architecture, we describe the remaining auxiliary components, all of which are implemented as Multi-Layer Perceptrons (MLPs). First, a 3D Projector aligns point cloud features to the same dimension, serving as supervise targets for the Latent CoT. Regarding the action generation mechanism, given that our fast acting expert adopts a Flow Matching policy (Black et al., 2024), specific modules are incorporated to handle the continuous generative dynamics. For the action expert input, we use a timestep MLP to encode the continuous time coordinate $t\in[0,1]$, initialized with sinusoidal embeddings, and a noised-action MLP to project the perturbed action state. For the action expert output, a projector MLP is used to transform the predicted flow velocity field.
3.3 Latent Spatio-Temporal Chain-of-Thought